Silver (I) oxide (Ag_2O) decomposes completely at
Answer #1

m = 1.6 g of Ag2O

V = 72.1 ml of O2

find mol of O2

n = (PV)/(RT) = 1*(72.1/1000)/(0.082*273) = 0.00322076297 mol of O2

then expect:

2Ag2O --> 4Ag + O2

ratio is 1:2

0.00322076297*2 = 0.00644152594 mol of Ag2O

mass = mol*MW = 0.00644152594*231.735 = 1.49272 g of Ag2O

mass of sample:

% = 1.49272/1.6 *100 = 93.295

a is the answer